OpenStudy (anonymous):

how do you solve for k? 6(2k-1/k)^2-22k+11/k=35

OpenStudy (cwrw238):

this expands to 24x^4 - 22k^3 - 59k^2 + 11k + 6 = 0 because of 24 and + 6 roots could be 2,4,6, 8 , 12 ..and the negatives of these f(2) = 0 so one zero = 2 dividing by (x - 2) to give a cubic expression would be the next step in the process there are another 3 roots to find

OpenStudy (cwrw238):

this is quite a laborious method but i can't figure out an easier way

OpenStudy (cwrw238):

graphing it would help

OpenStudy (anonymous):

I see. I tried this method and thought the same thing but it seems like the only way. Anyway, thank you! :)

OpenStudy (cwrw238):

yw

OpenStudy (cwrw238):

ah ! - a substitution can be made 6(2k-1/k)^2-22k+11/k=35 rearrange: 6(2k-1/k)^2-11(2k - 1/k) - 35 = 0 let 2k - 1 = x, then 6x^2 - 11x - 35 = 0 (3x + 5 )(2x - 7 ) = 0 x - -5/3 , 7/2 thus 2k - 1/k = 7/2 and 2k - 1/k = - 5/3 solve these to find the 4 values of k

OpenStudy (cwrw238):

how didn't i see that earlier!!???

OpenStudy (anonymous):

Thank you! :)

OpenStudy (cwrw238):

yw

OpenStudy (cwrw238):

its a lot easier this way....

OpenStudy (anonymous):

It is. :) I'll keep it in mind when something like this come up again.

OpenStudy (cwrw238):

yep - i should have seen it earlier - the 22k + 11/k and 2k - 1/k should have jogged my brain!

OpenStudy (cwrw238):

.. but maybe yesterday was a 'dog-day'

OpenStudy (cwrw238):

.. but maybe yesterday was a 'dog-day'